The data below show real wage levels at various percentiles of the wage distribution for
material-handling labourers in New Zealand

Percentile 1975 1995
10th 4.26 2.97
50th 6.02 3.89
90th 10.44 9.07

A. Calculate w90/w50, w50/w10, and w90/w10 percentile wage differentials for both
years.
B. Describe the changes in wage inequality in this local market
1
Expert's answer
2019-04-22T09:55:27-0400

A. In 1975: w90/w50 = 10.44/6.02 = 1.73,

w50/w10 = 1.41,

w90/w10 = 2.45.

In 1995: w90/w50 = 9.07/3.89 = 2.33,

w50/w10 = 1.31,

w90/w10 = 3.05.

B. The wage inequality increased in this local market.